# Dynamic Novel View Synthesis in High Dynamic Range

### v2

@@ 4.1 Quantitative evaluation
Competitors. We compare HDR-4DGS with latest alternatives: HexPlane (Cao & Johnson, 2023), 4DGS (Yang et al., 2024b), HDR-HexPlane (Wu et al., 2024a), HDR-NeRF (Huang et al., 2022) and HDR-GS (Cai et al., 2024). HDR-HexPlane preliminarily explores dynamic HDR synthesis by extending HexPlane with exposure conditioning and a static Sigmoid tone mapper. However, it lacks HDR supervision and omits explicit evaluation of HDR outputs, all of which have been addressed here for both fair comparison and completeness.
@@
Although HDR-GS and HDR-NeRF were originally designed for static scenes, we adapt and train them on dynamic scenes to underscore the importance of explicit spatial modeling in dynamic HDR reconstruction. All models are evaluated with their official implementations to ensure optimal performance.
@@
(I) Superior HDR DNVS quality. HDR-4DGS consistently outperforms all competing methods, including HDR-NeRF, HDR-GS, and HDR-HexPlane for HDR DNVS on HDR-4D-Syn, owing primarily to our explicit spatiotemporal HDR radiance modeling and the integration of a dynamic tone mapping mechanism that preserves temporal radiance coherence across views and exposures. In contrast, HDR-GS and HDR-NeRF rely on static tone mappers, failing to accurately bridge the HDR and LDR domains; while HDR-HexPlane achieves higher PSNR on HDR-4D-Real, we attribute this largely to the known limitations of PSNR as a metric—-particularly its tendency to favor overly smooth or blurred reconstructions (Li et al., 2020), as visually demonstrated in Fig. 7. As shown in Fig. 3 and videos in the supplementary material, HDR-4DGS preserves significantly sharper spatial structure and more accurate color details than all competitors.
Interestingly, it is observed from Tab. 3.4 that 4DGS with HDR supervision only achieves lower PSNR than HDR-4DGS with LDR supervision only. We attribute this to the extremely large dynamic range of HDR images, pixels in very bright regions dominate the gradients, while darker regions contribute little to the optimization. This imbalance causes the model to reconstruct bright areas more accurately but often fails to capture details in mid- and low-intensity regions, resulting in overly smooth or biased predictions.
@@ 4.3 Ablation study
To validate the effectiveness of the dynamic tone mapper, analyze its architectural design, and determine an appropriate temporal window length as a hyperparameter, we conduct ablation studies on the HDR-4D-Syn dataset.
Effect of dynamic tone mapper. To validate our dynamic tone mapper, we conduct comparative experiments against a plain MLP variant, two classical tone mappers Durand (Durand & Dorsey, 2000) and Reinhard (Reinhard et al., 2023) and the static tone mapper adopted by HDR-HexPlane. Additionally, ablation studies are performed to assess the contribution of extra pixel-level supervision (see Sec. 3.4). Key findings are: (I) Tab. 4.3 reveals that when replacing our dynamic tone mapper with other existing counterparts, we observe significant performance degradation. This confirms that explicit modeling of temporal radiance variations through our dynamic tone mapper is critical for maintaining HDR fidelity while existing mappers are inferior in doing that. Further, complementary ablation studies with respect to the scene representation verify that the observed performance gains are primarily attributable to our dynamic tone mapper, rather than 4DGS alone. (II) Disabling the pixel-level supervision leads to performance reduction (PSNR decreases by 1.03 dB), as shown in Tab. 4.3. This demonstrates that joint optimization with both ray-level and pixel-level constraints provides essential supervisory signals for learning physically plausible tone mapping operators.
Design of dynamic radiance context learner. The temporal radiance variation modeling constitutes a critical component of our dynamic tone mapper, as the temporal coherence of radiance decomposition directly influences HDR rendering performance. To systematically evaluate architectural suitability, we implement the dynamic radiance context learner with multiple sequence models and quantitatively compare their effectiveness. As shown in Tab. 4.3, while Transformer-based modules demonstrate theoretical advantages in long-term dependency modeling (Wen et al., 2023), our experiments reveal superior performance when employing GRU. This observation aligns with recent findings in sequence feature extraction, where GRU exhibits better efficiency in capturing local temporal patterns (Li et al., 2023).
@@
Durand (Durand & Dorsey, 2000) 22.85 0.825 0.195
@@
Temporal length. As described in Sec. 3.3, our dynamic tone mapper extracts radiance cues from the past $k$ timestamps in the radiance bank using a sliding window and a dynamic radiance context learner. The hyper-parameter $k$ governs the trade-off between temporal context coverage and computational efficiency: larger $k$ captures extended radiance dynamics but risk redundancy and noise sensitivity, while smaller $k$ prioritizes immediate temporal cues at the cost of modeling complex patterns. Through ablation experiments with $k=5,10,20,30$, as shown in Tab. 4.3, we find that $k=20$ achieves optimal performance, balancing responsiveness to dynamic scenes with practical efficiency constraints.
@@ 5 Conclusion
This work introduces the task of High Dynamic Range Dynamic Novel View Synthesis (HDR DNVS), addressing a critical limitation in prior HDR synthesis methods restricted to static scenes. We present HDR-4DGS, a novel framework designed to reconstruct temporally coherent HDR radiance fields and dynamic geometry from sparse, time-varying LDR observations. The key innovation of HDR-4DGS lies in its dynamic tone mapping module, which leverages a radiance bank and a dynamic radiance context learner to drive per-channel tone-mapping functions that adaptively bridge HDR and LDR domains across time. Extensive experiments demonstrate that HDR-4DGS significantly outperforms prior methods in HDR rendering fidelity, temporal radiance consistency, and computational efficiency. Limitations are discussed in Appendix A.3.
@@ A.1 Datasets
This section elaborates on the HDR-4D dataset introduced in this work. As summarized in Tab. 7, the dataset comprises three monocular acquisition configurations (Airplane, Deer, Lego and Tank) where distinct exposure times are applied across viewpoints, and four multi-view stereo configurations (Hook, Jump, Mutant and Standup) characterized by multi-exposure LDR images capture per viewpoint. This design enables comprehensive evaluation of dynamic scene reconstruction under varying lighting conditions and camera setups. Notably, we explicitly use different exposure times rather than exposure value to render multi-exposure images, which enables precise control over radiance sampling intervals, aligning with recent paper conventions. Fig. 5 and Fig. 6 show some example HDR images and LDR images with different exposure times of different scenes.
@@ A.2 Additional Experiments
Implementation details. Since Structure-from-Motion (SfM) (Schonberger & Frahm, 2016) struggles to perform reliably on dynamic monocular videos, we adopt different initialization strategies for synthetic and real-world scenes. Specifically, we randomly initialize $5\times 10^{4}$ Gaussians for each synthetic scene, while for real-world scenes, we initialize the Gaussians using a dense reconstructed point cloud and and the input images are downsampled by a factor of 4. In both cases, we maintain the same number of training iterations as 4DGS (Yang et al., 2024a). We conducted all the experiments on a single NVIDIA RTX 4090 GPU.
Additional visualization results. Fig. 8 and Fig. 9 present additional HDR DNVS rendering results on the HDR-4D-Syn and HDR-4D-Real datasets, respectively. Qualitative comparisons reveal that HDR-4DGS consistently preserves fine structural details and generates visually compelling renderings. In contrast, HDR-HexPlane tends to produce spatially blurred regions, while HDR-GS exhibits significant failure in reconstructing dynamic content, resulting in incomplete renderings.
@@
Furthermore, accurate rendering of continuous radiance distributions in HDR DNVS is essential for preserving high-fidelity radiance variations across complex lighting conditions, which often occur in the real-world dynamic scenes. As demonstrated in Fig. 12 and Fig. 13, our HDR-4DGS with dynamic tone mapper featured by a dynamic radiance context learner achieves superior temporal coherence in radiance transitions compared to baseline approaches, effectively capturing smoother details while maintaining photometric consistency. Meanwhile, Fig. 14 provides additional ablation study results of continuous radiance variation comparisons of HDR DNVS, which again, demonstrate the effectiveness of our method.

### v3

@@ 4.1 Quantitative evaluation
Competitors. We compare HDR-4DGS with latest alternatives: HexPlane (Cao and Johnson, 2023), 4DGS (Yang et al., 2024), HDR-HexPlane (Wu et al., 2024a), HDR-NeRF (Huang et al., 2022) and HDR-GS (Cai et al., 2024). HDR-HexPlane preliminarily explores dynamic HDR synthesis by extending HexPlane with exposure conditioning and a static Sigmoid tone mapper. However, it lacks HDR supervision and omits explicit evaluation of HDR outputs, all of which have been addressed here for both fair comparison and completeness.
@@
Although HDR-GS and HDR-NeRF were originally designed for static scenes, we adapt them on dynamic scenes to underscore the importance of explicit spatial modeling in dynamic HDR reconstruction. All models are tested with their official implementations to ensure optimal performance.
@@
(I) Superior HDR DNVS quality. HDR-4DGS consistently outperforms all competing methods, including HDR-NeRF, HDR-GS, and HDR-HexPlane for HDR DNVS on HDR-4D-Syn, owing primarily to our explicit spatiotemporal HDR radiance modeling and the integration of a dynamic tone mapping mechanism that preserves temporal radiance coherence across views and exposures. In contrast, HDR-GS and HDR-NeRF rely on static tone mappers, failing to accurately bridge the HDR and LDR domains; while HDR-HexPlane achieves higher PSNR on HDR-4D-Real, we attribute this largely to the noise in HDR ground truth and the known limitations of PSNR as a metric, particularly its tendency to favor overly smooth or blurred reconstructions (Li et al., 2020), as visually demonstrated in Fig. 9. As shown in Fig. 3 and videos in the supplementary material, HDR-4DGS preserves significantly sharper spatial structure and more accurate color details than all competitors. Although HDR-4DGS exhibits structural degradation in movable regions, we attribute this limitation to the inherent expressive capacity of our base representation model, 4DGS, rather than instability in HDR-4DGS’s handling of dynamic ranges. For further visual evidence, please refer to Fig. 10 and Fig. 11 in the Appendix where the vanilla 4DGS (Yang et al., 2024) suffers from even more severe structural distortion and geometry loss. In contrast, our HDR-4DGS alleviates these degradations to a noticeable extent, supporting the effectiveness of the proposed HDR modeling strategy.
Figure 4: Temporal variation with learned tone mapping patterns by DTM in two scenes.
@@
4.2 Dynamics in Tone Mapping Captured
To verify that our DTM learns dynamic rather than static tone mapping, we analyze scenes with pronounced motion-induced brightness changes (e.g., Airplane and Bed). After training HDR-4DGS, we extract the learned DTM module along with the radiance bank, both jointly define a dynamic mapping from HDR radiance to LDR counterpart, with condition on the time. To assess the dynamics of this learned mapping, we then sample different timestamps and visualize the corresponding tone-mapping curves over the HDR intensity, ass illustrated in Fig. 4.
We observe that (I) all curves are monotonically increasing (consistent with CRF definition (Huang et al., 2022)) yet exhibit scene-specific patterns, indicating adaptive tone reproduction; (II) In the Airplane scene, the red channel curve consistently lies above green and blue, reflecting the scene’s dominant reddish tone. As the airplane moves from shadows to brighter regions, curves shift upward over time, tracking increasing luminance; (III) In the Bed scene, curves gradually descend as the bed unfolds and ambient lighting dims. These temporal dynamics confirm DTM’s ability to adapt tone mapping to evolving lighting conditions and scene content.
@@ 4.4 Ablation study
We conduct ablation studies on the HDR-4D-Syn dataset (see Appendix A.2 for more experiments).
Effect of joint HDR reconstruction. To validate the effect of our DTM, we train 4DGS on single-exposure LDR images (exposure time 2.0s) and convert the synthesized LDR images of novel views to HDR images with a variety of existing Single HDR Image Reconstruction models: EIN (Liu et al., 2025b), IntrinsicHDR (Dille et al., 2024), KPNet (Wang et al., 2024a), KUNet (Wang et al., 2022).
Table 3: Results of 4DGS with independent HDR.
Method HDR
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
4DGS + EIN (Liu et al., 2025b) 17.75 0.770 0.185
4DGS + IntrinsicHDR (Dille et al., 2024) 11.63 0.673 0.251
4DGS + KPNet (Wang et al., 2024a) 20.92 0.813 0.197
4DGS + KUNet (Wang et al., 2022) 19.00 0.715 0.172
HDR-4DGS (Ours) 25.88 0.865 0.076
Tab. 3 shows that the two-stage pipeline significantly underperforms our HDR-4DGS since converting single-exposure LDR inputs to HDR is inherently ill-posed where the missing radiance information cannot be reliably recovered in isolation from the scene reconstruction process. Our integrated approach instead jointly optimizes radiance representation and novel-view synthesis, establishing its necessity and irreplaceability for high-fidelity HDR scene reconstruction.
Effect of dynamic tone mapper. We conduct comparative experiments against an MLP variant (part of HDR-GS (Cai et al., 2024) and HDR-NeRF (Huang et al., 2022), used for static tone-mapping), two classical tone mappers Durand (Durand and Dorsey, 2000) and Reinhard (Reinhard et al., 2023) and the static tone mapper adopted by HDR-HexPlane. Additionally, ablation studies are performed to assess the contribution of extra pixel-level supervision (see Sec. 3.4). Key findings are: (I) Tab. 4.4 reveals that when replacing our dynamic tone mapper with other existing counterparts, we observe significant performance degradation. This confirms that explicit modeling of temporal radiance variations through our dynamic tone mapper is critical for maintaining HDR fidelity while existing mappers are inferior in doing that. Further, complementary ablation studies with respect to the scene representation verify that the observed performance gains are primarily attributable to our dynamic tone mapper, rather than 4DGS alone. (II) Disabling the pixel-level supervision leads to performance reduction (PSNR decreases by 1.03 dB), as shown in Tab. 4.4. This demonstrates that joint optimization with both ray-level and pixel-level constraints provides essential supervisory signals for learning physically plausible tone mapping operators.
In addition, to rigorously evaluate the efficiency of our proposed DTM in preserving continuous radiance variations, we conduct ablation studies by substituting the DTM with an MLP baseline. Specifically, we synthesize temporally coherent HDR sequences for a dynamic scene using both architectures, as shown in Fig. 6 (also see Fig. 17) where our HDR-4DGS with DTM achieves superior temporal coherence in radiance transitions compared to baseline approach, validating the effectiveness of our DTM in preserving HDR coherence in dynamic domain.
@@
Durand (Durand and Dorsey, 2000) 22.85 0.825 0.195
@@
Temporal length. As described in Sec. 3.3, our proposed dynamic tone mapper (DTM) extracts radiance cues from the past $k$ timestamps in the radiance bank using a sliding window and a
Table 6: Analysis of the temporal context length.
Row $k$ HDR
@@
1 5 24.74 0.852 0.092
2 10 24.76 0.851 0.098
3 20 25.88 0.856 0.076
4 30 24.29 0.825 0.094
dynamic radiance context learner. The hyper-parameter $k$ governs the trade-off between temporal context coverage and computational efficiency: larger $k$ captures extended radiance dynamics but risk redundancy and noise sensitivity (e.g., scenes with large, rapid motions or complex illumination dynamics), while smaller $k$ prioritizes immediate temporal cues at the cost of modeling complex patterns (e.g., scenes characterized by minor motion or slow illumination variation). Through ablation experiments with $k$, as shown in Tab. 4.4, we find that $k=20$ achieves optimal performance, balancing responsiveness to dynamic scenes with practical efficiency constraints.
@@ 5 Conclusion
This work introduces High Dynamic Range Dynamic Novel View Synthesis (HDR DNVS), addressing a critical limitation in prior HDR synthesis restricted to static scenes. We present HDR-4DGS, a novel framework designed to reconstruct temporally coherent HDR radiance fields and dynamic geometry from sparse, time-varying LDR observations. The key of HDR-4DGS lies in its dynamic tone mapping module, which leverages a radiance bank and a dynamic radiance context learner to drive per-channel tone-mapping functions that adaptively bridge HDR and LDR domains across time. Extensive experiments demonstrate that HDR-4DGS significantly outperforms prior methods in HDR rendering fidelity, temporal radiance consistency, and computational efficiency.
@@ A.1 Datasets
This section elaborates on the HDR-4D-Syn and HDR-4D-Real datasets introduced in this work. As summarized in Tab. 7, the HDR-4D-Syn dataset is synthesized with Blender which comprises four monocular acquisition configurations (Airplane, Deer, Lego and Tank) where distinct exposure times are applied across viewpoints, and four multi-view stereo configurations (Hook, Jump, Mutant and Standup) characterized by multi-exposure LDR images capture per viewpoint. This design enables comprehensive evaluation of dynamic scene reconstruction under varying lighting conditions and camera setups. Notably, we explicitly use different exposure times rather than exposure value to render multi-exposure images, which enables precise control over radiance sampling intervals, aligning with recent paper conventions. Fig. 7 shows some example HDR images and LDR images with different exposure times of different scenes in HDR-4D-Syn.
@@
As for the HDR-4D-Real dataset, to ensure capture accuracy and stability, we employed a fixed multi-camera setup with six cameras securely mounted on tripods. Prior to introducing any dynamic elements, we captured multi-view static images of the empty scene. These static images were processed using COLMAP (Schonberger and Frahm, 2016) to compute precise intrinsic and extrinsic camera parameters, establishing a stable and accurate spatial reference frame. Crucially, all camera positions and optical parameters were rigorously maintained throughout the entire dynamic capture sequence, eliminating the need for recalibration and guaranteeing consistent geometric alignment across all timestamps. For each camera and each timestamp within the dynamic sequences, we captured three images with different exposure times (e.g., 1/120s, 1/50s, and 1/20s), enabling the subsequent synthesis of HDR ground truth. Finally, we adopt UltraFusion (Chen et al., 2025) to synthesize the corresponding HDR ground truth from multi-exposure LDR images. Tab. 8 summarizes the statistics of HDR-4D-Real, and Fig. 8 shows some example HDR images and LDR images with different exposure times of different scenes.
@@
Remark. It is noted that using identical synchronized cameras is indeed common practice in dynamic scene capture, as evidenced by major datasets like the Neural 3D Video Dataset (Li et al., 2022a) (18-21 synchronized cameras), Google Immersive Dataset (Broxton et al., 2020) (46-camera rig), and Technicolor Light Field Dataset (Sabater et al., 2017) (4×4 synchronized array). This approach is both a methodological convention and a practical necessity—synchronization ensures temporal consistency critical for dynamic reconstruction, while identical camera models eliminate radiometric calibration complexities from varying CRFs. Asynchronous capture or mixed camera setups would introduce significant additional challenges, including temporal misalignment artifacts, inconsistent motion blur, and complex cross-camera radiometric calibration that could obscure the core evaluation of HDR reconstruction capabilities. While robustness to such real-world variations represents valuable future work with extra challenges, this work focuses on common practice aligned with established conventions in the field.
@@ A.2 Additional Experiments
Implementation details. Since Structure-from-Motion (SfM) (Schonberger and Frahm, 2016) struggles to perform reliably on dynamic monocular videos, we adopt different initialization strategies for synthetic and real-world scenes. Specifically, we randomly initialize $5\times 10^{4}$ Gaussians for each synthetic scene, while for real-world scenes, we initialize the Gaussians using a dense reconstructed point cloud and and the input images are downsampled by a factor of 4. In both cases, we maintain the same number of training iterations as 4DGS (Yang et al., 2024). We conducted all the experiments on a single NVIDIA RTX 4090 GPU.
Ablation study of the design of dynamic radiance context learner. The temporal radiance variation modeling constitutes a critical component of our dynamic tone mapper, as the temporal coherence of radiance decomposition directly influences HDR rendering performance.
Table 9: Analysis of DRCL design.
Network HDR
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
RNN (Salehinejad et al., 2017) 25.63 0.847 0.100
LSTM (Staudemeyer and Morris, 2019) 25.53 0.845 0.106
GRU (Cho et al., 2014) 25.88 0.865 0.076
Transformer (Vaswani et al., 2017) 25.06 0.817 0.101
To systematically evaluate architectural suitability, we implement the dynamic radiance context learner with multiple sequence models and quantitatively compare their effectiveness. As shown in Tab. A.2, while Transformer-based modules demonstrate theoretical advantages in long-term dependency modeling (Wen et al., 2023), our experiments reveal superior performance when employing GRU. This observation aligns with recent findings in sequence feature extraction, where GRU exhibits better efficiency in capturing local temporal patterns (Li et al., 2023).
Additional visualization results. Fig. 10 and Fig. 11 provide additional visual comparisons of DNVS performance on the HDR-4D-Syn and HDR-4D-Real datasets. It is observed that 4DGS (Yang et al., 2024) consistently struggles to preserve fine-grained details and frequently exhibits structural degradation in movable regions. In contrast, our HDR-4DGS demonstrably retains superior detail fidelity, even when employing 4DGS as its underlying representation model. This outcome further corroborates the efficacy of our proposed DTM in enhancing detail preservation, as thoroughly discussed in Sec. 3.3.
Fig. 12 and Fig. 13 present additional HDR DNVS rendering results on the HDR-4D-Syn and HDR-4D-Real datasets, respectively. Qualitative comparisons reveal that HDR-4DGS consistently preserves fine structural details and generates visually compelling renderings. In contrast, HDR-HexPlane tends to produce spatially blurred regions, while HDR-GS exhibits significant failure in reconstructing dynamic content, resulting in incomplete renderings.
@@
Furthermore, accurate rendering of continuous radiance distributions in HDR DNVS is essential for preserving high-fidelity radiance variations across complex lighting conditions, which often occur in the real-world dynamic scenes. As demonstrated in Fig. 16 , our HDR-4DGS with dynamic tone mapper featured by a dynamic radiance context learner achieves superior temporal coherence in radiance transitions compared to baseline approaches, effectively capturing smoother details while maintaining photometric consistency. Meanwhile, Fig. 17 provides additional ablation study results of continuous radiance variation comparisons of HDR DNVS, which again, demonstrate the effectiveness of our method.
Figure 10: Additional visual comparison of DNVS on HDR-4D-Syn.
Figure 11: Additional visual comparison of DNVS on HDR-4D-Real.